OpenStudy (anonymous):

compute the area enclosed by the graphs of y = x + 6 , y = x^3 , and y = -x/2

OpenStudy (amistre64):

whats your graph of the bounded region ...

OpenStudy (anonymous):

sketched it out, and looks like left bounds are the intersection of the linear functions, and right bounds are the intersection of x + x and x^3

OpenStudy (anonymous):

so, -4 to 2

OpenStudy (anonymous):

just as an aside, how would you solve x + 6 = x^3

OpenStudy (amistre64):

complete the cube maybe, but i just wolfed it

OpenStudy (anonymous):

or it's obvious that it's two

OpenStudy (amistre64):

wasnt obvious to me till now lol

OpenStudy (anonymous):

or to me until I saw the wolf graph of the three

OpenStudy (amistre64):

ok, now is our boundary consistent, unchanging, along our interval?

OpenStudy (anonymous):

no, it changes at 0

OpenStudy (amistre64):

then what are your thoughts as to how we work this?

OpenStudy (anonymous):

calc from -4 to 0, add it to from 0 to 2

OpenStudy (amistre64):

very good, so whats our setup, what do you need help with

OpenStudy (anonymous):

reassurance, will post the final answer, these are much easier than some of the antiderivatives from previous chapters

OpenStudy (anonymous):

got 24 for A1

OpenStudy (amistre64):

\[\int_{-4}^{0}x+6+\frac12x~dx+\int_{0}^{2}x+6-x^3~dx\] \[\left.\frac12x^2+6x+\frac14x^2\right|_{-4}^{0}~+~\left. \frac12x^2+6x-\frac14x^4\right|_0^{2}\] \[\left.-\frac12(-4)^2-6(-4)-\frac14(-4)^2\right.~+~\left. \frac12(2)^2+6(2)-\frac14(2)^4\right.\]

OpenStudy (amistre64):

-8+24-4 = 12 for A1

OpenStudy (anonymous):

-3/4 * -64 - 24 = A1 ...is the area formula I cam up with after plugging in

OpenStudy (amistre64):

show me your antiderivative

OpenStudy (anonymous):

oops, cubed -4, should be squared

OpenStudy (anonymous):

got the same as you, (3/4)x^2 + 6x

OpenStudy (amistre64):

good :)

OpenStudy (anonymous):

0 - (3/4(-4)^2 - 24 = ...

OpenStudy (amistre64):

-6(-4) = +24

OpenStudy (anonymous):

oops, darn parenthesis

OpenStudy (anonymous):

12

OpenStudy (amistre64):

[3/4 (0)^2 + 6(0)] - [3/4 (-4)^2 +6(-4)] [0 + 0] - [12 -24] [0] - [-12] = 12

OpenStudy (anonymous):

A2 = 10

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

AA = 22

OpenStudy (amistre64):

very good

OpenStudy (anonymous):

woot, getting funner

OpenStudy (amistre64):

yeah, these are to build your confidence, theyll tear it down later ;)

OpenStudy (anonymous):

not me, probably done after this...bigger and better things to learn that computers can't do for us

OpenStudy (anonymous):

but yeah, this should be rudamentary for someone pursuing any sciences